# Neural cyberattacks applied to the vision under realistic visual stimuli

### v1

@@ 5.2 Impact study over the movie stimulus
Figure 9 shows the impact of FLO in the three attack instants selected. Figure 9(a) presents the number of spikes metric when attacking 25% of the neurons, where attacking during the gray event causes an increase of 12,297 spikes (78,12%). Then, when FLO was executed at the beginning of the movie (after the transition from the neutral event), the number of spikes had an increase of 43,394 spikes (60,57%). Finally, when FLO was executed in the middle of the movie, the attack caused an increase of 11,702 spikes (94,64%). After that, Figure  9(b) represents the results of the same three instants for 50% of the neurons. First, in the neutral event, the number of spikes metric augments in 29,422 spikes (186,92%) according to spontaneous activity. Secondly, in the instant corresponding to the beginning of the movie stimulus, there was an augment of 96,798 spikes (135,11%). Finally, during the movie, FLO causes an increase of 30,410 spikes (245,95%) compared to spontaneous behavior. In both number of neurons, it is clear that when the neurons have a lower response to an event, any overstimulation generates a notorious increase in spikes. However, if only the percentage of increase is considered, the most affected event is the instant during the movie. Nevertheless, from the number of spikes perspective, the greater impact can be observed during the sixth instant, representing the start of the movie stimulus, in both sets of neurons. This last situation can be explained since neurons change drastically from reacting to a neutral event to a new and more complex event, causing an increase of spikes in comparison to other events. In addition, the recovery to normal neuronal activity in the three attack instants is around 200-300 ms.
@@
Focusing on spike shifts, Figure 9(c) highlights that the gray event gets delayed by 70%, the second instant obtains a percentage of shifts of 75% and, during the last instant, the attack causes a variation of 93%. In contrast, Figure 9(d) represents the shifts percentage when 50% of the neurons are affected. The neutral event obtains 82% of shifts then, the event at the beginning of the movie gets delayed by 84% and, finally, during the movie stimulus, the percentage obtained is 98%. As in the last metric, the most sensitive effect corresponds to the movie stimulus since the neurons generate spikes in response to an event where they do not normally react too strongly. As an overview, all attack instants present a large shifts percentage that tends to decrease over time after the attack. Even at the end of the simulation, this metric maintains high values between 40%-50%.
@@
Figure 10 depicts the impact of JAM over the previously mentioned events. On the one hand, Figure 10(a) presents the number of spikes analysis when JAM inhibits 25% of the neurons. Attacking during the neutral event causes a decrease of 14,975 spikes (-95.14%), whereas performing JAM on the second event causes a decrement of 69,636 spikes (-97.20%). The interval corresponding to the middle of the movie stimulus falls to 11,714 spikes (-94.74%) when performing the attack. On the other hand, Figure 10(b) shows the same analysis as the previous figure for 50% of the neurons. The gray event suffers a decrease of 15,471 spikes (-98.29%), then JAM generates a reduction of 71,147 spikes (-99.31%) in the event representing the start of the movie, and the diminishment in the next event is 12,166 spikes (-98.40%). In this case, the neurons take seven intervals (700 ms) to return to spontaneous behavior when executing the attack in the gray event. In contrast, in the attack intervals corresponding to the beginning and the middle of the movie, the neurons take six intervals (600 ms) to recover to normal activity. As can be seen, both numbers of neurons have similar results; therefore, as in the previous stimulus, it is not necessary to attack many neurons to produce a notorious impact.
@@
In the analysis of the shifts, and particularly attending to Figure 10(c), the gray event gets 22% of shifts, while attacking the next event causes a variation of 66%. Finally, the percentage regarding the interval in the middle of the movie stimulus is 29%. Next, Figure 10(d) highlights that the shift percentages of neutral event is 7% when attacking 50% of the neurons. In contrast, when the attack was executed at the beginning of the movie, the shifts percentage was 44% and, finally, the shift percentage achieved when JAM was executed during the movie was 11%. The difference between sets of neurons is appreciable, unlike the previous metric. However, once the high peak passes, the attack intervals start decreasing until remaining in the range of 50%-60% for both numbers of neurons.
(a) Total number of spikes when FLO targets 25% of the neurons
(b) Number of spikes analysis for the impact over 50% of the neurons.
(c) Shifts percentages over 25% of the neurons
(d) Percentage of shifts when FLO is executed on 50% of the neurons.
@@
(a) Number of spikes of the impact over the 25% of the neurons.
(b) Total number of spikes when JAM affects the 50% of the neurons.
(c) Percentage of shifts of the attack impact, affecting the 25% of the neurons
(d) Shifts percentages when the attack is executed on 50% of the neurons.
@@ 5.3 Impact study over the drifting gratings stimulus
Figure 11 shows the evolution of FLO executed over the drifting gratings stimulus. Figure 11(a) indicates that FLO causes an augmentation of 9,673 spikes (90,43%) when executed on the neutral event according to attacking 25% of the neurons. The number of spikes when the attack is executed at the start of the drifting gratings event grows to 20,185 spikes (34,29%), even though attacking in the middle of the stimulus increases 20,899 more spikes (43,39%) than spontaneous activity. On the contrary, Figure 11(b) indicates that attacking during the gray event increases the activity by 24,959 spikes (233,35%), whereas applying FLO at the beginning of the stimulus makes the number of spikes increases by 53,337 spikes (90,61%). Finally, in the middle of the stimulus, the number of spikes increments by 52,466 spikes (108,92%). Additionally, the recovery in the three attack instants takes around two or three intervals (200-300 ms). In both numbers of neurons under attack, the gray event is the most affected, as in the previous stimuli, due to the neurons generating few spikes in their normal behavior, causing a significant impact in the base activity when an attack is executed. However, the remaining events when the attack is executed on the 25% of the neurons cause a lower number of spikes than the same events when FLO affected the 50% of the neurons. Moreover, the results highlight that, during the stimulus, the neural reaction to the attack is greater compared to visualizing a neutral stimulus.
@@
Regarding the shift percentage metric, Figure 11(c) indicates that the impact during the first instant is 69% of shifts delayed. The value when the attack was executed at the beginning of the stimulus is 95%, and, finally, the percentage obtained when FLO was used during the drifting gratings is 94% of variation. In contrast, Figure 11(d) shows that the neutral event obtains 84% of shifts; then, the instant corresponding to the beginning of the stimulus gets delayed by 98%. Lastly, the impact obtained in the middle of the stimulus is 97%. These results highlight high percentages in the whole simulation, remaining over 80% except for the instant of the gray event when the attack affects 25% of the neurons. Unlike the number of spikes metric where the neurons, after the attack, return to their normal activity, the shifts analysis shows that there is a great delay in each interval. Therefore, although the neurons generate a similar number of spikes as the spontaneous behavior, it does not mean that there is not a substantial delay in time.
@@
Finally, Figure 12 presents the impact of JAM when it attacks the 25% of the neurons and the 50% of the neurons. On the one hand, Figure 12(a) highlights that JAM decreases the number of spikes by 10,361 spikes (-96.87%) in the gray event, while, in the second attack instant, the number of spikes has 56,387 spikes less (-95.79%). In contrast, the diminishment obtained when JAM was used during the drifting gratings is 45,119 spikes (-93.67%). On the other hand, Figure 12(b) depicts that, for the neutral event, the number of spikes decreases by 10,613 spikes (-99.22%). Then, the results corresponding to the beginning of the event highlight 58,031 spikes less (-98.58%), although the attack instant during the stimulus decrements by 47,121 spikes (-97.83%). Besides, the neurons attacked in the instants corresponding to the gray event and in the middle of the stimulus return to their spontaneous activity after nine intervals (900 ms), while the neurons attacked at the beginning of drifting gratings require eight intervals (800 ms) to return to their normal behavior. In this case, the attack affects both sets of neurons equally, highlighting the low complexity to make a significant impact.
@@
The percentage of shifts analysis is also divided by the number of neurons attacked. Figure 12(c) indicates that the impact obtained in the neutral event is 44%. Then, the percentage when the attack was executed at the beginning and during the drifting gratings stimulus is the same, 31%. In contrast, Figure 12(d) represents that attacking the gray event causes a delay of 17%, whereas the following instants have the same percentage of shifts, 11%. The results in the attack intervals, especially within the drifting gratings stimulus, suggest that the delay is the same regardless of when the attack is executed, where both attack instants have a similar evolution over the simulation. Nevertheless, in the remaining intervals, the percentage remains stable, over 80%.
@@
(a) Total number of spikes when the attack is executed on 25% of the neurons.
(b) Number of spikes of the impact over the 50% of the neurons.
(c) Shift analysis of FLO when is executed on 25% of the neurons.
(d) Percentage of shifts of attack impact, affecting the 50% of the neurons.
@@
(a) Number of spikes analysis of JAM when is executed on 25% of the neurons.
(b) Total number of spikes of the impact over the 50% of the neurons.
(c) Shifts percentages of JAM impact executed on 25% of the neurons.
(d) Percentages of shifts when the attack is executed on 50% of the neurons.
@@ 6 Discussion
This section compares different relevant aspects of the results presented in Section 5, such as the impact of neural cyberattacks on different visual stimuli, the difference between the cyberattacks according to the impact metrics, and the number of neurons attacked. Besides, this section compares these results with those existing in the literature.
@@ 6.2 Impact comparison between cyberattacks
It is necessary to analyze the impact of both cyberattacks individually to compare them. Focusing on FLO, and attending to the percentage of spikes increase compared to spontaneous behavior, the results indicate that the impact of FLO is highly dependent on the stimuli, getting a minimum percentage of increase of 12,74% and a maximum of 94,64% when 25% of the neurons are attacked. In contrast, when FLO is used over 50% of the neurons, the minimum percentage of increase is 32,83%, and the maximum is 245,95%. As can be observed, the minimum and maximum percentages of both sets of neurons coincide with the same events, being the OFF-flash the event with the lowest percentage and the middle of the movie stimulus the one with the highest percentage. In contrast, the overstimulation caused in terms of the number of spikes generated by FLO oscillates between 9,673 and 43,394 spikes for 25% of the neurons, while in the 50%, FLO oscillates between 24,959 and 96,798 spikes. These outcomes demonstrate that this attack fluctuated according to the event attacked. Secondly, the experimentation for JAM proves that this threat can cause a considerable damage regardless of the stimulus, causing a reduction of spikes between 93% and 99% compared to spontaneous signaling. Regarding the propagation of the effect caused by each attack, FLO propagates until 200-300 ms (2-3 intervals) depending on the stimuli evaluated, while JAM gets a propagation duration between 600 ms and 1,100 ms (6-11 intervals). In addition, the shifts analysis presents that both cyberattacks produced a substantial delay in the attack instants within the drifting gratings stimulus. Furthermore, the results depict that both attacks induce a delay of the neurons over the whole simulation, as can be seen in the shifts percentage figures, although the neurons seem to return to the spontaneous activity in the number of spikes metric.

### v2

@@ 5.2 Impact study over the movie stimulus
Figure 9 shows the impact of FLO in the three attack instants selected. The left plot of Figure 9(a) presents the number of spikes metric when attacking 25% of the neurons, where attacking during the gray event causes an increase of 12,297 spikes (78.12%). Then, when FLO was executed at the beginning of the movie (after the transition from the neutral event), the number of spikes had an increase of 43,394 spikes (60.57%). Finally, when FLO was executed in the middle of the movie, the attack caused an increase of 11,702 spikes (94.64%). After that, the right figure represents the results of the same three instants for 50% of the neurons. First, in the neutral event, the number of spikes metric augments in 29,422 spikes (186.92%) according to spontaneous activity. Secondly, in the instant corresponding to the beginning of the movie stimulus, there was an augment of 96,798 spikes (135.11%). Finally, during the movie, FLO causes an increase of 30,410 spikes (245.95%) compared to spontaneous behavior. In both number of neurons, it is clear that when the neurons have a lower response to an event, any overstimulation generates a notorious increase in spikes. However, if only the percentage of increase is considered, the most affected event is the instant during the movie. Nevertheless, from the number of spikes perspective, the greater impact can be observed during the sixth instant, representing the start of the movie stimulus, in both sets of neurons. This last situation can be explained since neurons change drastically from reacting to a neutral event to a new and more complex event, causing an increase of spikes in comparison to other events. In addition, the recovery to normal neuronal activity in the three attack instants is around 200-300 ms.
@@
Focusing on spike shifts, Figure 9(b) highlights, in the plot that represents 25% of the neurons attacked, that the gray event gets delayed by 70%, the second instant obtains a percentage of shifts of 75% and, during the last instant, the attack causes a variation of 93%. In contrast, the right side of Figure 9(b) represents the shifts percentage when 50% of the neurons are affected. The neutral event obtains 82% of shifts then, the event at the beginning of the movie gets delayed by 84% and, finally, during the movie stimulus, the percentage obtained is 98%. As in the last metric, the most sensitive effect corresponds to the movie stimulus since the neurons generate spikes in response to an event where they do not normally react too strongly. As an overview, all attack instants present a large shifts percentage that tends to decrease over time after the attack. Even at the end of the simulation, this metric maintains high values between 40%-50%.
@@
Figure 10 depicts the impact of JAM over the previously mentioned events. On the one hand, the left plot of Figure 10(a) presents the number of spikes analysis when JAM inhibits 25% of the neurons. Attacking during the neutral event causes a decrease of 14,975 spikes (-95.14%), whereas performing JAM on the second event causes a decrement of 69,636 spikes (-97.20%). The interval corresponding to the middle of the movie stimulus falls to 11,714 spikes (-94.74%) when performing the attack. On the other hand, the plot that represents 50% of the neurons attacked on Figure 10(a) shows the same analysis as the previous figure for 50% of the neurons. The gray event suffers a decrease of 15,471 spikes (-98.29%), then JAM generates a reduction of 71,147 spikes (-99.31%) in the event representing the start of the movie, and the diminishment in the next event is 12,166 spikes (-98.40%). In this case, the neurons take seven intervals (700 ms) to return to spontaneous behavior when executing the attack in the gray event. In contrast, in the attack intervals corresponding to the beginning and the middle of the movie, the neurons take six intervals (600 ms) to recover to normal activity. As can be seen, both numbers of neurons have similar results; therefore, as in the previous stimulus, it is not necessary to attack many neurons to produce a notorious impact.
@@
In the analysis of the shifts, and particularly attending to the left plot of Figure 10(b), the gray event gets 22% of shifts, while attacking the next event causes a variation of 66%. Finally, the percentage regarding the interval in the middle of the movie stimulus is 29%. Next, the right figure highlights that the shift percentages of neutral event is 7% when attacking 50% of the neurons. In contrast, when the attack was executed at the beginning of the movie, the shifts percentage was 44% and, finally, the shift percentage achieved when JAM was executed during the movie was 11%. The difference between sets of neurons is appreciable, unlike the previous metric. However, once the high peak passes, the attack intervals start decreasing until remaining in the range of 50%-60% for both numbers of neurons.
(a) Total number of spikes when FLO targets 25% (left image) and 50% (right image) of the neurons. Events under attacks are indicated by red numbers and gray shades.
(b) Shifts percentages over 25% (left image) and 50% (right image) of the neurons. The red numbers indicated the three events under attack individually.
@@
(a) Number of spikes of the impact over the 25% (left image) and 50% (right image) of the neurons. Attack instants are highlighted by red numbers.
(b) Percentage of shifts of the attack impact, affecting 25% (left image) and 50% (right image) of the neurons. Events under attack are highlighted using red numbers and gray shading.
@@ 5.3 Impact study over the drifting gratings stimulus
Figure 11 shows the evolution of FLO executed over the drifting gratings stimulus. The left plot of Figure 11(a) indicates that FLO causes an augmentation of 9,673 spikes (90.43%) when executed on the neutral event according to attacking 25% of the neurons. The number of spikes when the attack is executed at the start of the drifting gratings event grows to 20,185 spikes (34.29%), even though attacking in the middle of the stimulus increases 20,899 more spikes (43.39%) than spontaneous activity. On the contrary, the figure that represents the 50% of the neurons in Figure 11(a) indicates that attacking during the gray event increases the activity by 24,959 spikes (233.35%), whereas applying FLO at the beginning of the stimulus makes the number of spikes increases by 53,337 spikes (90.61%). Finally, in the middle of the stimulus, the number of spikes increments by 52,466 spikes (108.92%). Additionally, the recovery in the three attack instants takes around two or three intervals (200-300 ms). In both numbers of neurons under attack, the gray event is the most affected, as in the previous stimuli, due to the neurons generating few spikes in their normal behavior, causing a significant impact in the base activity when an attack is executed. However, the remaining events when the attack is executed on the 25% of the neurons cause a lower number of spikes than the same events when FLO affected the 50% of the neurons. Moreover, the results highlight that, during the stimulus, the neural reaction to the attack is greater compared to visualizing a neutral stimulus.
@@
Regarding the shift percentage metric, the plot showing 25% of the neurons depicted in Figure 11(b) indicates that the impact during the first instant is 69% of shifts delayed. The value when the attack was executed at the beginning of the stimulus is 95%, and, finally, the percentage obtained when FLO was used during the drifting gratings is 94% of variation. In contrast, the right figure shows that the neutral event obtains 84% of shifts; then, the instant corresponding to the beginning of the stimulus gets delayed by 98%. Lastly, the impact obtained in the middle of the stimulus is 97%. These results highlight high percentages in the whole simulation, remaining over 80% except for the instant of the gray event when the attack affects 25% of the neurons. Unlike the number of spikes metric where the neurons, after the attack, return to their normal activity, the shifts analysis shows that there is a great delay in each interval. Therefore, although the neurons generate a similar number of spikes as the spontaneous behavior, it does not mean that there is not a substantial delay in time.
@@
Finally, Figure 12 presents the impact of JAM when it attacks the 25% of the neurons and the 50% of the neurons. On the one hand, the plot corresponding to 25% of the neurons from Figure 12(a) highlights that JAM decreases the number of spikes by 10,361 spikes (-96.87%) in the gray event, while, in the second attack instant, the number of spikes has 56,387 spikes less (-95.79%). In contrast, the diminishment obtained when JAM was used during the drifting gratings is 45,119 spikes (-93.67%). On the other hand, the figure illustrating 50% of the neurons presented in Figure 12(b) depicts that, for the neutral event, the number of spikes decreases by 10,613 spikes (-99.22%). Then, the results corresponding to the beginning of the event highlight 58,031 spikes less (-98.58%), although the attack instant during the stimulus decrements by 47,121 spikes (-97.83%). Besides, the neurons attacked in the instants corresponding to the gray event and in the middle of the stimulus return to their spontaneous activity after nine intervals (900 ms), while the neurons attacked at the beginning of drifting gratings require eight intervals (800 ms) to return to their normal behavior. In this case, the attack affects both sets of neurons equally, highlighting the low complexity to make a significant impact.
@@
The percentage of shifts analysis is also divided by the number of neurons attacked. The left side of Figure 12(b) indicates that the impact obtained in the neutral event is 44%. Then, the percentage when the attack was executed at the beginning and during the drifting gratings stimulus is the same, 31%. In contrast, the plot corresponding to 50% of the neurons from Figure 12(b) represents that attacking the gray event causes a delay of 17%, whereas the following instants have the same percentage of shifts, 11%. The results in the attack intervals, especially within the drifting gratings stimulus, suggest that the delay is the same regardless of when the attack is executed, where both attack instants have a similar evolution over the simulation. Nevertheless, in the remaining intervals, the percentage remains stable, over 80%.
@@
(a) Total number of spikes when the attack is executed on 25% (left plot) and 50% (right plot) of the neurons. Impacts under attack are emphasized using red numbers and gray shades.
(b) Shift analysis of FLO when is executed on 25% (left plot) and 50% (right plot) of the neurons. Red numbers and gray shades marked the events affected by the attack.
@@
(a) Number of spikes analysis of JAM when it is executed on 25% (left plot) and 50% (right plot) of the neurons. Red digits and gray tones remark on instants affected by attack.
(b) Percentages of shifts when the attack is executed on 25% (left plot) and 50% (right plot) of the neurons. Red indicators and gray-scale overlays highlighted the events under attack.
@@ 6 Discussion
This section compares different relevant aspects of the results presented in Section 5, such as the impact of neural cyberattacks on different visual stimuli, the difference between the cyberattacks according to the impact metrics, and the number of neurons attacked. Besides, this section compares these results with those existing in the literature. Furthermore, it presents an analysis either the effects on the brain, the mitigation strategies, and the scalability of neural cyberattacks.
@@ 6.2 Impact comparison between cyberattacks
It is necessary to analyze the impact of both cyberattacks individually to compare them. Focusing on FLO, and attending to the percentage of spikes increase compared to spontaneous behavior, the results indicate that the impact of FLO is highly dependent on the stimuli, getting a minimum percentage of increase of 12.74% and a maximum of 94.64% when 25% of the neurons are attacked. In contrast, when FLO is used over 50% of the neurons, the minimum percentage of increase is 32.83%, and the maximum is 245.95%. As can be observed, the minimum and maximum percentages of both sets of neurons coincide with the same events, being the OFF-flash the event with the lowest percentage and the middle of the movie stimulus the one with the highest percentage. In contrast, the overstimulation caused in terms of the number of spikes generated by FLO oscillates between 9,673 and 43,394 spikes for 25% of the neurons, while in the 50%, FLO oscillates between 24,959 and 96,798 spikes. These outcomes demonstrate that this attack fluctuated according to the event attacked. Secondly, the experimentation for JAM proves that this threat can cause a considerable damage regardless of the stimulus, causing a reduction of spikes between 93% and 99% compared to spontaneous signaling. Regarding the propagation of the effect caused by each attack, FLO propagates until 200-300 ms (2-3 intervals) depending on the stimuli evaluated, while JAM gets a propagation duration between 600 ms and 1,100 ms (6-11 intervals). In addition, the shifts analysis presents that both cyberattacks produced a substantial delay in the attack instants within the drifting gratings stimulus. Furthermore, the results depict that both attacks induce a delay of the neurons over the whole simulation, as can be seen in the shifts percentage figures, although the neurons seem to return to the spontaneous activity in the number of spikes metric.
@@ 6.4 Comparison between the literature and the current research
6.5 Analysis of the effects on the brain and mitigation strategies.
The execution path for neural cyberattacks in the real world begins with the exploitation of vulnerabilities existing in a neurostimulation device. In this regard, López Bernal et al. [9] identified several cybersecurity flaws in emerging neurostimulation implants, including those used in Neuralink. These vulnerabilities enable an attacker to gain control over the device and manipulate its operation. The consequences of unauthorized access include alterations in normal brain activity, which can result in adverse effects on the patient. In this context, the literature has documented that inadequate or malicious neurostimulation can lead to severe physical and psychological consequences in the brain. This risk is particularly critical in devices such as BCIs, where the artificial generation or alteration of brain signals may result in serious health risks, as described by López Bernal et al. [8]. For instance, during neurostimulation therapies such as those used for Parkinson’s disease, cyberattacks could induce the inability to initiate voluntary movement, memory difficulties, or general weakness as neurological side effects. Psychological consequences may include anxiety, confusion, or depression. In extreme cases, these attacks could even provoke death. In the context of visual neurostimulation, potential effects might include temporary or permanent blindness, inability to perceive motion, or inability to recognize objects
Focusing on possible protection countermeasures to detect neural cyberattacks and mitigate their effects, the following mechanisms could be applied. Firstly, as a detection technique, artificial intelligence can be used to identify anomalies by identifying unusual patterns or significant deviations in neural activity. These AI-based systems enhance the capability to detect subtle changes that may indicate the presence of an attack. Additionally, implementing real-time alert systems, whether directed at human operators or automated control systems, aims to facilitate rapid intervention in the event of detected threats. Furthermore, logging mechanisms can be integrated to systematically record spike activity, input stimuli, external perturbations, and timing information for the detection of attacks. Secondly, within mitigation measures, the logging technique, as mentioned previously, provides a foundation for post-attack analysis, enabling traceability, correlation between attack events and their effects, and support for the development of future mitigation strategies based on the actions taken in response to the attacks. Besides, generating reports on the effects and scope of neural cyberattacks would contribute to minimizing their impact and improving the security of BCIs. To interrupt the propagation of the attack, temporarily stopping BCI operation can prevent the escalation of the attack impact before it compromises the patient’s health. Finally, conducting a session to educate both clinicians and patients about the formative sessions could help reduce the risks of neurostimulation BCIs and their consequences, as humans are usually the weakest element in security systems.
6.6 Evaluating the Scalability of Neural Cyberattacks
In both real and simulation scenarios, there is direct access to individual neurons, either through multiple electrodes in the case of BCIs or through the neuronal topology used in the simulation. Therefore, the methodology used to execute the attacks, applied in this case to V1, can be extended to other brain regions, such as the motor or prefrontal cortex, since the approach is generalizable. However, changing the target of the attack introduces several new challenges. In real BCIs, the intruder must understand the functional characteristics of the neurons in the new cortical region, including their connectivity patterns, firing dynamics, and general behavior. Each brain area has distinct structural and functional properties, which complicates the design and deployment of effective attacks. In simulation scenarios, the attack code must be adapted to the new realistic neural topology once the corresponding region has been analyzed and modeled. This adjustment includes reconfiguring the input of the model, adjusting attack parameters such as the number of targeted neurons or the attack interval, and validating the results. In both real and simulated contexts, it becomes necessary to evaluate how the new area responds to cyberattacks, analyze the specific vulnerabilities it may present, and perform comparative studies between results.